OpenStudy (anonymous):

∫1/(15√(2pi))*e^-(x-268)²/450

OpenStudy (anonymous):

you will not find a closed form for the "anti derivative" of \[e^{-(x-268)^2}\]

OpenStudy (anonymous):

\[\int\limits_{250}^{280}1/15\sqrt{2\pi}\times e ^-{\left(x-268\right)^{2}/450 dx}\]

OpenStudy (anonymous):

It's tricky to type this up with the equation formulas available. Anyway, it's basically a probability problem with a mean and SD.

OpenStudy (anonymous):

yes so i think you are going to have to use a table of some kind because you are not going to find the anti derivative of this beast

OpenStudy (anonymous):

Problem is the lecturer seems to expect us to be able to integrate this because no tables have been recommended or given. Brand new to probabilities.

OpenStudy (zarkon):

There is no antiderivative in terms of elementary functions so you will need to use a table or calculator/computer to compute this integral

OpenStudy (anonymous):

Ok, the lecturer gave us these problems over the weekend without much instruction, so how do I use a calculator to solve this? Do I plug in the limits to get a numeric value and then integrate everything? Doesn't seem correct.

OpenStudy (zarkon):

what type of calculator do you have?

OpenStudy (anonymous):

Just a plain scientific one

OpenStudy (zarkon):

does it have a numerical integration function on it?

OpenStudy (anonymous):

I doubt it; I don't see any integral symbols. What should I look for?

OpenStudy (zarkon):

i get .788144601395 on my calculator

OpenStudy (zarkon):

\[\int\limits_{50}^{280}\frac{1}{15\sqrt{2\pi}}e^{-\frac{(x-268)^2}{450}}dx\approx.788144601395\]

OpenStudy (anonymous):

That sounds correct, given that the question asked, "What percentage of pregnancies are considered full term?" 79% sounds about right. How did you get that answer? What do I use on my calculator?

OpenStudy (zarkon):

I used my Ti-nspire CX CAS

OpenStudy (anonymous):

What? Ok, definitely don't have those symbols on my calculator. Not sure why the lecturer gave us this problem, but thanks for your help, anyway.

OpenStudy (zarkon):

I have 3 scientific calculators that can do numerical integrals...what type of scientific calculator do you have?

OpenStudy (anonymous):

Oh, sorry, you were referring to the actual name of the calculator. Mine is TI-30XA

OpenStudy (zarkon):

ti36x pro casio fx-115es sharp EL-W516x all can do numerical integrals

OpenStudy (zarkon):

the ti-30 can't do numerical integral

OpenStudy (zarkon):

I have a ti-30xs

OpenStudy (zarkon):

I am a calculator whore...I have 11 or 12 of them

OpenStudy (anonymous):

Alright, that solves that problem. Provided I know that one needs a calculator in order to solve this, I'm relieved. (I like your comment about your affinity for calculators :-))